OpenStudy (anonymous):

The perimeter of a rectangle is 24ft. the length is 2ft longer than the width find the demensions. write a system of linear equations and solve the system, X is the length, Y width. 2x+2x= x=y what is the length and what is the width in feet?

OpenStudy (anonymous):

x + 2 = length x = width x + 2 + x = 24 2x + 2 = 24 2x = 22 x = 11 The width is 11ft, and since the length is two feet longer than the width, the length is 13ft.

OpenStudy (ganpat):

given, X = length and Y = width also, X = Y + 2 perimenter for rectangle, P = 2 (X +Y) Therefore, 24 = 2 (Y + 2 + Y) 24 = 4Y + 4 y = 5 width Y = 5 Length X =7

OpenStudy (anonymous):

totally confused. There are two different answers

OpenStudy (ganpat):

ohh.. @JBrightman3 : Perimeter for rectangles shd be 2 (l + B), right ??

OpenStudy (anonymous):

I'm sorry, my answer was wrong. I only counted one side of the length and one side of the width. I should have counted two of each. @Ganpat has the correct answer. I apologize again!

OpenStudy (anonymous):

It's okay, thanks for the help. so if I write the first equation would it be 2(x+y)? and what about the second equation

OpenStudy (anonymous):

Yes, it would be 2(x + y), but you don't know what y is. You just know that y is x + 2. So it is now 2(x + x + 2) = 24. Solve for x. 2(2x + 2) = 24 4x + 4 = 24 4x = 20 x = 5 So the width is 5, and the length is 2 feet longer than the width, so the length is 7.

OpenStudy (anonymous):

so its not 13 length and 11 width? it is width 5 and length 7?

OpenStudy (anonymous):

Exactly. :)

OpenStudy (anonymous):

x=intercept2 and y intercept 2/5/ what is the slope of the line?

OpenStudy (anonymous):

The y intercept is 2 / 5? If that is so, then the equation of the line is 2y + 0.4x = 0.8. Put that into slope-intercept form.

OpenStudy (anonymous):

2y + 0.4x = 0.8 y + 0.2x = 0.4 y = -0.2x + 0.4 The slope intercept form is y = mx + b, where m is the slope. In this equation, -0.2 is m, so -0.2 is the slope. :)

OpenStudy (anonymous):

@antjojo3009 Is that one of the options?